You buy a commemorative coin for $20. The value of the coin increases at a rate of 3.5% per year.
How much will the coin be worth in 15 years??
After how many years will the coin have doubled in value??

Answer by Alan3354(69443)

You buy a commemorative coin for $20. The value of the coin increases at a rate of 3.5% per year.
How much will the coin be worth in 15 years??
After how many years will the coin have doubled in value??
===============
You didn't spec a compounding period, so I'll use years.
If it's 3.5%/12 per month, the result would be different.
----
FV = PV*(1 + r)^t --- FV = future value, PV = present value
FV = 20*(1 + 0.035)^15
FV =~ $33.507
----
A gain of $13.50, a BFD
-------------
After how many years will the coin have doubled in value??
FV = 2*PV*(1.o35)^t
40 = 20*1.035^t
2 = 1.035^t
ln(2) = t*ln(1.035)
t = ln(2)/ln(1.035)
t =~ 20.15 years